# Oka-1 manifolds: New examples and properties

### v2

@@ 4. Simply connected algebraically elliptic manifolds are aOka-1
Rational manifolds and algebraically elliptic manifolds form two subclasses of the class of unirational manifolds. It has recently been shown by Kaliman and Zaidenberg [KalimanZaidenberg2024X] that every smooth cubic hypersurface in $\mathbb{P}^{n}$, $n>2$, is algebraically elliptic. This seems to give the first examples of non-rational projective manifolds which are algebraically elliptic. On the other hand, it is unknown whether every rational manifold is algebraically elliptic.
@@ 5. Condition LSAP and its functorial properties
The sprays in Definition 5.2 can either be defined over unspecified open neighbourhoods of the compact disc $D$, or else they could be continuous over $D$ and holomorphic over its interior $\mathring{D}$. The application of this condition in the proofs of Propositions 5.4 and 5.5 works in both cases, since the technique of gluing sprays on Cartan pairs works in both cases (see [12, Sect. 5.9]). As noted in [2, Remark 7.12], LSAP is equivalent to the ostensibly weaker condition that every dominating holomorphic spray $F:D\times\mathbb{B}^{N}\to V$ can be approximated by sprays on a given bigger disc $D^{\prime}\supset D$ as in Definition 5.2. Furthermore, it suffices to ask that this condition hold with $N=\dim X$. In fact, LSAP is a localisation of the convex approximation property $\mathrm{CAP}_{1+\dim X}$ (see [12, Definition 5.4.3]).
@@
Pick a point $x\in X$. By the hypothesis there are a holomorphic map $h:Y\to X$ and a point $y\in Y$ such that $h(y)=x$, $h$ is submersive at $y$, and $Y$ satisfies LSAP at $y$. Let $U_{0}\subset Y$ be a neighbourhood of $y$ such that the condition in Definition 5.2 holds for sprays of discs in $Y$ with values in $U_{0}$. Since $h$ is a submersion at $y$, the implicit function theorem gives a smaller neighbourhood $U\subset U_{0}$ of $y$ such that the restriction $h:U\to V:=h(U)\subset X$ is fibrewise isomorphic to the projection $U\cong V\times W\to V$, where $W$ is a neighbourhood of $y$ in the fibre $Z=h^{-1}(x)$. In this local trivialization we write $y=(x,z)$ with $z\in Z$. Let $D\subset D^{\prime}$ be a special pair of discs (see Definition 5.1) and $F:D\times\mathbb{B}^{N}\to V$ be a holomorphic spray. Let $G=(F,z):D\times\mathbb{B}^{N}\to U\cong V\times W$ be the lifting of $F$ with the constant second component $z$. By the assumption, there is a number $10<r<$ 1 such that $G$ can be approximated on $D\times r\mathbb{B}^{N}$ by holomorphic maps $\widetilde{G}:D^{\prime}\times r\mathbb{B}^{N}\to Y$. The holomorphic map $\widetilde{F}=h\circ\widetilde{G}:D^{\prime}\times r\mathbb{B}^{N}\to X$ then approximates $F$ on $D\times r\mathbb{B}^{N}$. This shows that $X$ has LSAP at the point $x$. Since $x$ was arbitrary, $X$ has LSAP.
@@
Acknowledgements. Forstnerič is supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291, J1-3005, and N1-0237 from ARIS, Republic of Slovenia. A part of the work on this paper was done during his visit to the University of Adelaide in February 2024, and he wishes to thank this institution for hospitality. The authors wish to thank Erlend F. Wold for consultations which led to the proof of Proposition 2.4.

### v3

@@ 4. Simply connected algebraically elliptic manifolds are aOka-1
Rational manifolds and algebraically elliptic manifolds form two subclasses of the class of unirational manifolds. It has recently been shown by Kaliman and Zaidenberg [25] that every smooth cubic hypersurface in $\mathbb{P}^{n}$, $n>2$, is algebraically elliptic. This seems to give the first examples of non-rational projective manifolds which are algebraically elliptic. On the other hand, it is unknown whether every rational manifold is algebraically elliptic.
@@ 5. Condition LSAP and its functorial properties
The sprays in Definition 5.2 can either be defined over unspecified open neighbourhoods of the compact disc $D$, or else they could be continuous over $D$ and holomorphic over its interior $\mathring{D}$. The application of this condition in the proofs of Propositions 5.4 and 5.5 works in both cases, since the technique of gluing sprays on Cartan pairs works in both cases (see [12, Sect. 5.9]). As noted in [2, Remark 7.12], LSAP is equivalent to the ostensibly weaker condition that every dominating holomorphic spray $F:D\times\mathbb{B}^{N}\to V$ can be approximated by sprays on a given bigger disc $D^{\prime}\supset D$ as in Definition 5.2.
@@
Pick a point $x\in X$. By the hypothesis there are a holomorphic map $h:Y\to X$ and a point $y\in Y$ such that $h(y)=x$, $h$ is submersive at $y$, and $Y$ satisfies LSAP at $y$. Let $U_{0}\subset Y$ be a neighbourhood of $y$ such that the condition in Definition 5.2 holds for sprays of discs in $Y$ with values in $U_{0}$. Since $h$ is a submersion at $y$, the implicit function theorem gives a smaller neighbourhood $U\subset U_{0}$ of $y$ such that the restriction $h:U\to V:=h(U)\subset X$ is isomorphic to the projection $U\cong V\times W\to V$ by a fibre-preserving map, where $W$ is a neighbourhood of $y$ in the fibre $Z=h^{-1}(x)$. In this local trivialization we write $y=(x,z)$ with $z\in Z$. Let $D\subset D^{\prime}$ be a special pair of discs (see Definition 5.1) and $F:D\times\mathbb{B}^{N}\to V$ be a holomorphic spray. Let $G=(F,z):D\times\mathbb{B}^{N}\to U\cong V\times W$ be the lifting of $F$ with the constant second component $z$. By the assumption, there is a number $10<r<$ 1 such that $G$ can be approximated on $D\times r\mathbb{B}^{N}$ by holomorphic maps $\widetilde{G}:D^{\prime}\times r\mathbb{B}^{N}\to Y$. The holomorphic map $\widetilde{F}=h\circ\widetilde{G}:D^{\prime}\times r\mathbb{B}^{N}\to X$ then approximates $F$ on $D\times r\mathbb{B}^{N}$. This shows that $X$ has LSAP at the point $x$. Since $x$ was arbitrary, $X$ has LSAP.
@@
Acknowledgements. Forstnerič is supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291, J1-3005, and N1-0237 from ARIS, Republic of Slovenia. A part of the work on this paper was done during his visit to the University of Adelaide in February 2024, and he wishes to thank this institution for hospitality. The authors wish to thank Erlend F. Wold for consultations which led to the proof of Proposition 2.4. They also thank two referees for carefully reading the paper and suggesting several improvements.